OpenStudy (anonymous):

How do I solve a question like this: x/5=7/4

OpenStudy (anonymous):

cross multiplication multiply left hand side and right hand side, both with 5 x=35/4

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

medal, plz xD

OpenStudy (anonymous):

you cross multiply it. so, you'll get 4x=5(7) 4x=35 x=35/4